# DISCODE: Distribution-Aware Score Decoder for Robust Automatic Evaluation of Image Captioning

### v1

@@ B.4 Results on FOIL
Table 13 compares DISCODE with the baselines on the FOIL benchmark (Shekhar et al. 2017), which is specifically designed to assess a model’s ability to detect hallucinated content in image captions. As shown, DISCODE outperforms all baselines, including reference-based ones, demonstrating strong hallucination detection capabilities and further validating its robustness as a reference-free evaluation metric.
@@
| RefPAC-S (Sarto et al. 2023) | 93.7 | 94.9 |
| Polos (Wada et al. 2024) | 93.3 | 95.4 |
| PAC-S (Sarto et al. 2023) | 89.9 | 89.9 |
| FLEUR (Lee, Park, and Kang 2024) | 96.8 | 96.8 |
| G-VEval (Tong et al. 2025) | 97.8 | 98.4 |
@@ B.5 Full Results
Tables 17 and 17 show the full results corresponding to Table 6, which compares performance across four LLMs: LLama-3-8B (Dubey et al. 2024), Vicuna-13B (Chiang et al. 2023), Nous-Hermes-2-Yi-34B (Research 2024), and Qwen-1.5-72B (Team 2024). As shown, performance improvements are consistent across all domains in MCEval, as well as both Kendall’s and in Flickr8k-Expert, Flickr8k-CF and Composite.
@@
Tables 18 and 19 shows the full results corresponding to Table 7 with six state-of-the-art LLMs: InternVL-2.5-8B, InternVL-2.5-78B (Chen et al. 2024b, a), Qwen2-VL-7B-Instruct, Qwen2-VL-72B-Instruct (Bai et al. 2023; Wang et al. 2024a), CogVLM2-Chat-19B (Wang et al. 2024b; Hong et al. 2024), Mini-CPM-V-2.6 (Hu et al. 2024; Yao et al. 2024). Consistent with the results for LLMs, we observed performance improvements across both Kendall’s and and most subcategories in Pascal-50S.
@@ C Conventional Benchmarks
This dataset includes 3995 images, combining 997 images from Flickr8k, 991 images from Flickr30k (Young et al. 2014), and 2,007 images from MSCOCO (Lin et al. 2015). Each image is paired with three captions, and human evaluators rated each caption’s relevance to the corresponding image on a scale from 1 to 5.

### v2

@@ Acknowledgments
| Metric | FF | Flickr8k-Expert | Flickr8k-CF | Composite | Pascal-50S | ||||||||||
| Kendall | Kendall | Kendall | Kendall | Kendall | Kendall | HC | HI | HM | MM | Mean | |||||
| BLEU-1 (Papineni2002BLEU) | – | 32.2 | 32.3 | 17.9 | 9.3 | 29.0 | 31.3 | 64.6 | 95.2 | 91.2 | 60.7 | 77.9 | |||
| BLEU-4 (Papineni2002BLEU) | – | 30.6 | 30.8 | 16.9 | 8.7 | 28.3 | 30.6 | 60.3 | 93.1 | 85.7 | 57.0 | 74.0 | |||
| ROUGE (Lin2004ROUGE) | – | 32.1 | 32.3 | 19.9 | 10.3 | 30.0 | 32.4 | 63.9 | 95.0 | 92.3 | 60.9 | 78.0 | |||
| METEOR (Banerjee2005METEOR) | – | 41.5 | 41.8 | 22.2 | 11.5 | 36.0 | 38.9 | 66.0 | 97.7 | 94.0 | 66.6 | 81.1 | |||
| CIDEr (Vedantam et al. 2015) | – | 43.6 | 43.9 | 24.6 | 12.7 | 34.9 | 37.7 | 66.5 | 97.9 | 90.7 | 65.2 | 80.1 | |||
| SPICE (Anderson2016SPICE) | – | 51.7 | 44.9 | 24.4 | 12.0 | 38.8 | 40.3 | 63.6 | 96.3 | 86.7 | 68.3 | 78.7 | |||
| BERT-S (Zhan2020BERTScore) | – | 39.2 | 22.8 | – | – | 30.1 | 65.4 | 96.2 | 93.3 | 61.4 | 79.1 | ||||
| LEIC (Cui2018LEIC) | 46.6 | – | 29.5 | – | – | – | – | – | – | – | – | ||||
| BERT-S++ (Yi2020BERTScorepp) | – | 46.7 | – | – | – | 44.9 | 65.4 | 98.1 | 96.4 | 60.3 | 80.1 | ||||
| TIGEr (Jiang2019TIGEr) | – | 49.3 | – | – | – | 45.4 | 56.0 | 99.8 | 92.8 | 74.2 | 80.7 | ||||
| ViLBERTScore (Lee2020ViLBERTScore) | – | 50.1 | – | – | – | 52.4 | 49.9 | 99.6 | 93.1 | 75.8 | 79.6 | ||||
| MID (Kim2022MID) | – | 54.9 | 37.3 | – | – | – | 67.0 | 99.7 | 97.4 | 76.8 | 85.2 | ||||
| CLAIR (Chan2023CLAIR) | – | 48.3 | 38.2 | – | – | 61.0 | 52.4 | 99.5 | 89.8 | 73.0 | 78.7 | ||||
| RefCLIP-S (Hessel2021CLIPS) | 52.6 | 53.0 | 36.4 | 18.8 | 51.2 | 55.4 | 64.9 | 99.5 | 95.5 | 73.3 | 83.3 | ||||
| RefPAC-S (Sarto2023PACS) | 55.5 | 55.9 | 37.6 | 19.5 | 53.0 | 57.3 | 67.7 | 99.6 | 96.0 | 75.6 | 84.7 | ||||
| RefPAC-S++ViT-B (Sarto2024PACScorepp) | 55.3 | 55.7 | 37.9 | 19.6 | 54.7 | 59.1 | 67.2 | 99.6 | 96.2 | 74.2 | 84.5 | ||||
| RefPAC-S++ViT-L (Sarto2024PACScorepp) | – | 57.9 | 38.8 | – | – | 61.6 | – | – | – | – | 84.7 | ||||
| Polos (Wada2024PolosPolaris) | – | 56.4 | 37.8 | – | – | 57.6 | 70.0 | 99.6 | 97.4 | 79.0 | 86.5 | ||||
| DENEB (ViT-L) (Matsuda et al. 2024) | – | 56.8 | 38.3 | – | 58.2 | 76.1 | 99.7 | 97.4 | 77.9 | 87.8 | |||||
| RefFLEUR (Lee et al 2024)) | – | 51.9 | 38.8 | – | – | 64.2 | 68.0 | 99.8 | 98.0 | 76.1 | 85.5 | ||||
| G-VEval-C (GPT-4o) (Tong2025GVEval) | 60.5 | 58.7 | 38.2 | 19.9 | – | – | – | – | – | – | – | ||||
| RefDISCODE-LV (Ours) | 56.1 | 56.5 | 40.4 | 20.9 | 61.0 | 66.0 | 70.6 | 99.8 | 98.6 | 82.1 | 87.8 | ||||
| RefDISCODE-IN (Ours) | 57.1 | 57.5 | 40.5 | 20.9 | 60.5 | 64.9 | 68.0 | 99.7 | 97.7 | 79.7 | 86.3 | ||||
| Method | Flickr8k-Expert | Flickr8k-CF | Composite | Pascal-50S | |||||||||||
Kendall |
Kendall |
Kendall |
Kendall |
Kendall |
Kendall |
HC |
HI |
HM |
MM |
Mean |
| DISCODE | 55.7 | 56.1 | 40.2 | 20.8 | 61.1 | 66.0 | 62.9 | 99.8 | 97.9 | 77.4 | 84.5 | ||||
| w/o cross-entropy term | 54.3 | 54.6 | 39.5 | 20.4 | 58.3 | 63.1 | 62.5 | 99.8 | 97.2 | 75.7 | 83.8 | ||||
| w/o divergence term | 49.6 | 49.9 | 39.9 | 20.6 | 59.6 | 64.4 | 61.7 | 99.7 | 97.7 | 72.7 | 83.0 | ||||
| w/o adaptive definition for | 55.2 | 55.6 | 40.2 | 20.8 | 60.5 | 65.4 | 63.9 | 99.8 | 98.1 | 75.5 | 84.3 | ||||
| Variance | Flickr8k-Expert | Flickr8k-CF | Composite | Pascal-50S | |||||||||||
Kendall |
Kendall |
Kendall |
Kendall |
Kendall |
Kendall |
HC |
HI |
HM |
MM |
Mean |
| 1.0 | 55.6 | 55.9 | 40.2 | 20.8 | 60.9 | 65.9 | 61.8 | 99.8 | 97.9 | 77.2 | 84.2 | ||||
| 0.5 | 55.6 | 55.9 | 40.2 | 20.8 | 60.9 | 65.9 | 61.8 | 99.8 | 97.9 | 77.2 | 84.2 | ||||
| 0.1 | 55.7 | 56.1 | 40.2 | 20.8 | 61.1 | 66.0 | 62.9 | 99.8 | 97.9 | 77.4 | 84.5 | ||||
| 0.05 | 55.7 | 56.1 | 40.2 | 20.8 | 61.1 | 66.0 | 62.9 | 99.7 | 97.9 | 77.4 | 84.5 | ||||
| 0.01 | 55.7 | 56.1 | 40.2 | 20.8 | 61.1 | 66.0 | 63.0 | 99.7 | 97.9 | 76.5 | 84.3 | ||||
| Method | LLM | Backbone (sec) | Scoring (sec) | Overhead (%) |
| Raw score | LLama-3-8B | 0.00 | ||
| FLEUR | LLama-3-8B | 0.23 | ||
| DISCODE w/o Analytical Solution | LLama-3-8B | 9.46 | ||
| DISCODE | LLama-3-8B | 0.73 | ||
| Raw score | Vicuna-13B | 0.00 | ||
| FLEUR | Vicuna-13B | 0.24 | ||
| DISCODE w/o Analytical Solution | Vicuna-13B | 11.5 | ||
| DISCODE | Vicuna-13B | 0.79 | ||
| Raw score | Hermes-Yi-34B | 0.00 | ||
| FLEUR | Hermes-Yi-34B | 0.18 | ||
| DISCODE w/o Analytical Solution | Hermes-Yi-34B | 8.13 | ||
| DISCODE | Hermes-Yi-34B | 0.66 | ||
| Raw score | Qwen-1.5-72B | 0.00 | ||
| FLEUR | Qwen-1.5-72B | 0.04 | ||
| DISCODE w/o Analytical Solution | Qwen-1.5-72B | 1.90 | ||
| DISCODE (w/ AS) | Qwen-1.5-72B | 0.11 |
@@ B.4 Results on FOIL
Table 13 compares DISCODE with the baselines on the FOIL benchmark (shekhar-etal-2017-foil), which is specifically designed to assess a model’s ability to detect hallucinated content in image captions. As shown, DISCODE outperforms all baselines, including reference-based ones, demonstrating strong hallucination detection capabilities and further validating its robustness as a reference-free evaluation metric.
@@
| RefPAC-S (Sarto2023PACS) | 93.7 | 94.9 |
| Polos (Wada2024PolosPolaris) | 93.3 | 95.4 |
| PAC-S (Sarto2023PACS) | 89.9 | 89.9 |
| FLEUR (Lee2024FLEUR) | 96.8 | 96.8 |
| G-VEval (Tong2025GVEval) | 97.8 | 98.4 |
@@ B.5 Full Results
Tables 17 and 17 show the full results corresponding to Table 6, which compares performance across four LLMs: LLama-3-8B (Dubey2024Llama3), Vicuna-13B (vicuna2023), Nous-Hermes-2-Yi-34B (NousResearch2024HermesYi), and Qwen-1.5-72B (QwenTeam2024qwen1.5). As shown, performance improvements are consistent across all domains in MCEval, as well as both Kendall’s and in Flickr8k-Expert, Flickr8k-CF and Composite.
@@
Tables 18 and 19 shows the full results corresponding to Table 7 with six state-of-the-art LLMs: InternVL-2.5-8B, InternVL-2.5-78B (Chen2024InternVL; Chen2024InternVL2_5), Qwen2-VL-7B-Instruct, Qwen2-VL-72B-Instruct (Bai2023QwenVL; Wang2024Qwen2VL), CogVLM2-Chat-19B (Wang2023CogVLM; Hong2024CogVLM2), Mini-CPM-V-2.6 (Hu2024MiniCPM; Yao2024MiniCPMv). Consistent with the results for LLMs, we observed performance improvements across both Kendall’s and and most subcategories in Pascal-50S.
@@ C Conventional Benchmarks
This dataset includes 3995 images, combining 997 images from Flickr8k, 991 images from Flickr30k (young-etal-2014-image), and 2,007 images from MSCOCO (lin2015microsoftcococommonobjects). Each image is paired with three captions, and human evaluators rated each caption’s relevance to the corresponding image on a scale from 1 to 5.